Construct a h********** linear differential equation with constant coefficients whose solution is y = 6+3xe^x–cosx


1
Expert's answer
2022-02-15T17:25:20-0500

Let us construct a homogeneous linear differential equation with constant coefficients whose solution is "y = 6+3xe^x–\cos x."

It follows that the characteristic equation of this differential equation should be the following:

"k(k-1)^2(k^2+1)=0." Indeed, in this case the general solution is "y=C_1+(C_2+C_3x)e^x+C_4\cos x+C_5\sin x," and hence we get the solution "y = 6+3xe^x–\cos x" when "C_1=6,\ C_2=0,\ C_3=3,\ C_4=-1,\ C_5=0."

The characteristic equation is equvalent to

"(k^3-2k^2+k)(k^2+1)=0,"

and hence to

"k^5-2k^4+2k^3-2k^2+k=0."

Therefore, a homogeneous linear differential equation with constant coefficients whose solution is "y = 6+3xe^x–\cos x" is the following:

"y^{(V)}-2y^{(IV)}+2y^{(III)}-2y''+y'=0."